Find values for m and b in the following system so that the solution to te system is ( -3, 4).
5x + 7y = b
mx + y = 22

Answer by Earlsdon(6294)

Find the values of m and b given the solution to the system is (-3, 4):
5x + 7y = b
mx +y = 22
Substitute the x- and y-coordinates from the given point (-3, 4) into each equation and solve for the b and m.
5(-3) + 7(4) = b Simplify.
-15 + 28 = b
b = 13
mx + y = 22
m(-3) + 4 = 22 Subtract 4 from both sides of the equation.
-3m = 18 Divide both sides by -3.
m = -6
You can verify these solutions by substituting b = 13 and m = -6 into the appropriate equation and doing the math.
Try this yourself.